A 10.0 kg block is pulled by a 20.0N force at an angle of 37 with the horizontal. If the friction between the block and the surface is negligible, the acceleration of the block is:

A. 1,20 m/s2

B. 120m/s2

C. 160m/s2

D. 16,0m/s2

E. 1,60m/s2

1
Expert's answer
2021-01-19T07:11:28-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_x=ma_x=ma,""Fcos\theta=ma,""a=\dfrac{Fcos\theta}{m},""a=\dfrac{20.0\ N\cdot cos37^{\circ}}{10.0\ kg}=1.60\ \dfrac{m}{s^2}."

Answer:

E) "a=1.60\ \dfrac{m}{s^2}."